05/27/99 owl shift summary

Sandrine Laplace (laplace@hep.uchicago.edu)
Thu, 27 May 1999 06:56:12 -0500

Shift Summary for owl shift, 05/27/97

On Shift:
Sandrine Laplace
Hogan Nguyen

other people were present until 1 or 2 am : RSK, SachaG, JasonH

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7

Accounting of lost time:

-- "Unkown crate error" due to a bad DYC in crate 24. Only the 4th and
last TDC header words in this crate are corrupted when ever we get the
message. Swapted the DYC, it seems to have fixed the problem.

-- Search for DYC BUSY too long durations (>12 microsec.) for 1 or 2
hours.
It could explain why there are 2 different readout deadtimes:
11.3 microsec. no L2
13.3 microsec. HCC in trigger
It seems that they are 5 FERA ADC DYC BUSY signals (4616's nu.26 :
2_01, 2_02, 2_03, 2_04, 2_05) which occasionnaly raise duration greater
than 12 microsec.
We checked the events sizes of these crates, they indeed are up to
500 bytes occasionnaly. Could this be due to incorrect pedestals ?
I let to Rick the task to give the conclusion.

-- Run number 13147 screwed up while measuring the DYC BUSY durations.
I don't know if it is due to a bad manipulation while triggering
with stop/start buttons, or if it screwed up without my help...
We stopped the run, but the stop session screwed up, too (I asked
Hogan to spell me the word "screw up", I realized how helpfull it will
be).
A fresh start sequence restablished the order.

-- Ed updated the reference for the laser gain drift plot.

- HN worked on hot wires in DC1 and DC2. The edge wires were masked.
Reduced DA noise pickup on DC1. DC4 occasionally gets a noise burst.
HN added ores to cables going to TDC. It seems ok now.

- Started new random pulser run 13150.

========================================================================
Runs ended this shift:
Run 13147 - RANDOM PULSER (screwed up at the end)
Run 13148 - RANDOM PULSER
Run 13149 - LASER SCAN
Run 13150 - RANDOM PULSER (still running at the end of the shift)

Tapes brought to FCC this shift: 0

========================================================================
Accesses during shift: opened access

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

********************************************************************
* LAPLACE Sandrine *
* --------------- *
* Enrico Fermi Institute KTeV 5514,S Blackstone Apt 318*
* 5640, S Ellis ave. at Chicago, IL 60637 *
* Chicago, IL 60637, USA Fermilab USA *
* http://www.geocities.com/Nashville/Rodeo/1190 *
********************************************************************